# Berkovich Motives

### v1

@@ 6. Free motivic sheaves
We already know that all truncations of $M\otimes\mathbb{Q}$ and $M\otimes\mathbb{Q}/\mathbb{Z}$ remain in $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$. To conclude, it suffices to control the transition map. Thus, take any $N,N^{\prime}\in\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$ that are concentrated in degree 00 and such that $N$ is rational while $N^{\prime}$ is torsion. We need to see that the image of any map $f:N\to N^{\prime}$ is in $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$. As this image is necessarily torsion, we have to see invariance under change of algebraically closed field. Thus let $C\subset C^{\prime}$ be an inclusion of non-discrete algebraically closed Banach fields over $X$. We want to see that the image of $N(C)\to N^{\prime}(C)$ maps isomorphically to the image of $N(C^{\prime})\to N^{\prime}(C^{\prime})$, noting that we know $N^{\prime}(C)=N^{\prime}(C^{\prime})$. Thus, the map is necessarily injective. But if a class $\alpha\in N^{\prime}(C)$ lies in the image of $N(C^{\prime})\to N^{\prime}(C^{\prime})=N^{\prime}(C)$, then by finitaryness there is some Banach-$C$-algebra $R$ of topologically finite type such that the image of $\alpha$ under $N^{\prime}(C)\to N^{\prime}(R)$ lies in the image of $N(R)\to N^{\prime}(R)$. But then there is a section $R\to C$, showing that $\alpha$ is already in the image of $N(C)$. ∎
@@ 7. The Cancellation Theorem
$-\otimes\mathbb{Z}(1):\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)\to\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$
@@
By arc-descent, we can assume that $X$ is strictly totally disconnected. As $\mathbb{Z}(1)$ is a compact object of $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$ in that case, one can reduce further to the case of algebraically closed Banach fields, and only the nonarchimedean case is nontrivial. Also, the result is clear with torsion coefficients, so we will work with $\mathbb{Q}$-coefficients below.
@@
splitting the previous displayed map on $\pi_{i}$. Actually, we will rather have a map from the source to eventually defined sequences in the target, i.e. the colimit over all $N$ of $\mathrm{Map}(\mathbb{N}_{\geq N},-)$. Still, it will be a splitting in the obvious sense.
The left-hand side can be computed as the $i$-th homotopy group of the rationalization of the animated commutative monoid
@@
for some $k,\ell\geq 1$; and two such maps yield the same element if, possibly after refinement and addition of some section $\mathbb{T}\to\mathrm{Sym}^{k^{\prime}}(U\times(\overline{\mathbb{G}}_{m},\ast))$, they admit a homotopy from $\mathbb{T}\times\mathrm{sd}^{\ell^{\prime}}\Delta^{i+1}_{C}\times\mathbb{B}$.
@@
This finishes the construction. It is easy to check that for all $N$ for which the map is defined, it is a splitting. Indeed, in that case $p_{1}=p_{2}$ and so $p_{1}^{N}=p_{2}$ cuts out the open balls around the $N-1$-th roots of unity (counted with multiplicity in case $N-1$ is divisible by the residue characteristic of $C$). The map to $\mathbb{N}[U]$ is constant, so $Y_{1}$ yields $N$ times the original section of $\mathbb{N}[U]$, while $Y_{2}$ yields it $N-1$ times; the difference is thus the original section. This already proves injectivity of the map
@@
[…]
@@
Of course, once this lemma was used to finish the proof of the cancellation theorem, we see that $\mathrm{End}(\mathbb{Z}(n))=\mathrm{End}(\mathbb{Z})$ which is concentrated in degree 00 and hence the triviality of the action of each $\sigma$ implies that the action of $\Sigma_{n}$ on $\mathbb{Z}(n)$ is homotopy-coherently trivial. For the moment, the weaker assertion is sufficient.
@@
There is an action of $\mathrm{GL}_{2}$ on $\mathbb{A}^{2}\setminus\{0\}$. We want to see that the antidiagonal matrix, switching the two coordinates, acts trivially on the reduced homology. But the subgroup generated by unipotent matrices acts trivially, by homotopy-invariance. This subgroup is $\mathrm{SL}_{2}$. Up the action of $\mathrm{SL}_{2}$, the antidiagonal matrix is equivalent to the diagonal matrix with entries $(-1,1)$. It remains to see that multiplication by $-1$ (or any element of $\mathbb{G}_{m}$) acts trivially on the reduced homology of $\mathbb{G}_{m}$, which is a direct verification using the identification with $\overline{\mathbb{G}}_{m}$. Indeed, the reduced homology of $\mathbb{G}_{m}$ is the kernel of $\mathbb{Z}_{\mathrm{mot}}[\mathbb{G}_{m}]\to\mathbb{Z}_{\mathrm{mot}}[\ast]$, and the isomorphism with $\overline{\mathbb{G}}_{m}$ is given by sending $\sum n_{i}[x_{i}]$ to $\prod\overline{x}_{i}^{n_{i}}$. The action of $-1$ sends $\sum n_{i}[x_{i}]$ to $\sum n_{i}[-x_{i}]$ which is mapped to $(-1)^{\sum n_{i}}\prod\overline{x}_{i}^{n_{i}}$. But $\sum n_{i}=0$ for any element in the kernel of $\mathbb{Z}_{\mathrm{mot}}[\mathbb{G}_{m}]\to\mathbb{Z}_{\mathrm{mot}}[\ast]$ and so $(-1)^{\sum n_{i}}=1$. ∎
@@ 8. $K$-theory
Of course, these results are also true with $\mathbb{Z}$-coefficients.
@@
can be shown to be $\mathbb{Z}(n)[2n]$ as follows. Use the covering $\mathbb{P}^{n}$ by $\mathbb{A}^{n}$ and $\mathbb{P}^{n}\setminus\{0\}$; the first piece is contractible, the second piece is an $\mathbb{A}^{1}$-fibration over $\mathbb{P}^{n-1}$ and hence cancels with $\mathbb{P}^{n-1}$, so the remainder is a shift of the reduced homology of the intersection $\mathbb{A}^{n}\setminus\{0\}$, which gives $\mathbb{Z}(n)[2n]$.) ∎
@@
Here, departing from our usual conventions, we consider a functor with values in spectra, not $\mathcal{D}(\mathbb{Z})$. As we will be working rationally momentarily, the difference will disappear. Note however that everything we did with $\mathcal{D}(\mathbb{Z})$-coefficients works also with $\mathrm{Sp}$-coefficients.
@@
Explicitly, $\mathbb{S}[\ast/\mathbb{G}_{m}]$ is the group completion of
@@
indeed, the displayed object is the free $E_{\infty}$-monoid on $\ast/\mathbb{G}_{m}$. We may then compute the rationalization of the group completion in terms of the Breen–Deligne resolution involving tensor products of terms
@@
embedding the normalizer of the diagonal torus, and it suffices to see that they induce isomorphisms
@@
We want to make $\overline{K}$ more explicit. For this, we first have to replace $K_{\mathrm{cn}}$ by a finitary approximation. We assume that our base ring $A$ is equipped with a unit $T\in A$ with $|T|<1$ and such that $T$ admits $n$-th roots for infinitely many $n$. In that case the non-unital ring $A_{<1}\subset A$ is Tor-unital: If $A_{<1}$ is an ideal in a unital ring $B$, then it is automatically the colimit of $B$ along multiplication by $T^{1/n_{i}-1/n_{i+1}}$ (where $n_{i}$ is a strictly increasing sequence of integers for which $T$ admits $n_{i}$-th roots) and hence flat and idempotent. This implies that the fiber of $K(B)\to K(B/A_{<1})$ is independent of $B$ (it is the $K$-theory of the dualizable category of almost modules, and even that dualizable category is independent of $B$); we denote it by $K(A_{<1})$. Let $K_{\mathrm{cn}}(A_{<1})$ denote its connective cover. There is a natural map $K_{\mathrm{cn}}(A_{<1})\to K_{\mathrm{cn}}(A)$.
@@
The functor $A\mapsto\mathrm{cofib}(K_{\mathrm{cn}}(A_{<1})\to K_{\mathrm{cn}}(A))$, as well as its nonconnective version $\mathrm{cofib}(K(A_{<1})\to K(A))$, commutes with filtered colimits of Banach rings. When evaluated on totally disconnected $A$, the connective and nonconnective versions agree.
@@
as $A=(\mathbb{Z}\oplus A_{<1})[T^{-1}]$. Passing to the cofiber of $K(A_{<1})$ the middle term becomes simply $K(\mathbb{Z})$ which is certainly invariant under changing $A$; and the left-most term is unchanged under replacing $A_{<1}$ by its $T$-adic completion, by the results of Thomason–Trobaugh [TT90]. This gives the result for nonconnective $K$-theory. The case of connective $K$-theory follows.
@@
After inverting all primes $p$ such that $|p|<1$ somewhere on $\mathcal{M}(A)$, the functor $A\mapsto\mathrm{cofib}(K(A_{<1})\to K(A))$ on strictly totally disconnected $A$ defines a finitary arc-sheaf of spectra that is ball-invariant, and hence agrees with $\overline{K}$.
@@
After inverting the residue characteristic $p$ of $C$, this is a direct consequence of the identification with $\overline{K}(C)$. But after $p$-completion, $K(\mathcal{O}_{C})$ and $K(C)$ agree, so this cofiber agrees with $K(k)$ after $p$-completion, and this is just $\mathbb{Z}_{p}$ as $k$ is a perfect field. This has a unique $E_{\infty}$-$K(C)$-algebra structure, which necessarily glues to the ring structure after inverting $p$. ∎
@@
We can work with the fiber instead. This is the $K$-theory of the dualizable stable $\infty$-category $\mathcal{D}_{\mathrm{tor}}(A_{\leq 1})^{a}$ of torsion almost modules over $A_{\leq 1}$. Indeed, there is a Verdier sequence
@@
But $\mathcal{D}_{\mathrm{tor}}(A_{\leq 1})^{a}$ localizes on $\mathcal{M}(A)$ by the usual properties of the structure sheaf of perfectoid algebras, which in particular yields a sheaf of idempotent algebras with values in $\mathcal{D}_{\mathrm{tor}}(A_{\leq 1})^{a}$ (when interpreting torsion modules as complete modules). Then also the $K$-theory localizes, as desired. ∎
@@
On perfectoid Banach rings $A$ (with $|p|_{A}<1$) such that all sheaves on the compact Hausdorff space $\mathcal{M}(A)$ are hypersheaves, the functor $A\mapsto\mathrm{cofib}(K(A_{<1})\to K(A))_{\mathbb{Q}}$ is a finitary ball-invariant hypercomplete arc-sheaf. In particular, on such $A$, we have
@@
Rationally, finite étale descent holds because of trace maps; it even gives pro-finite étale descent. This combines with the analytic descent we already have to give a hypercomplete arc-sheaf. Finally, we check invariance under perfectoid balls (which is equivalent to usual ball-invariance). We know explicitly the value on $C\langle T^{1/p^{\infty}}\rangle_{1}$ (as $\mathcal{M}(C\langle T^{1/p^{\infty}}\rangle_{1})$ is a tree on which all sheaves are hypersheaves): This is the cofiber of
@@
In particular, restricting to $A$ such that all sheaves on the compact Hausdorff space $\mathcal{M}(A)$ are hypersheaves, the functor
@@
For the first statement, we can use the finitary nature of the functor to do a noetherian approximation to assume that $A_{\leq 1}$ is noetherian. Then this follows from Cisinski’s cdh-descent for $KH$, [Cis13]. Alternatively, in order to give an argument independent of the existing motivic formalism, it follows from pro-cdh descent [Mor16, Theorem 3.7] (for the equal characteristic 00 case that we really need) and [KST18, Theorem A] (in the general case) upon making things ball-invariant. Namely, pro-cdh descent implies that the pro-spectrum $\{\mathrm{cofib}(K(A_{<r})\to K(A))\}_{r}$ satisfies descent, but after making it ball-invariant, all terms agree. Here $K(A_{<r})$ is notation for the fiber of $K(A_{\leq 1})\to K(A_{\leq 1}/A_{<r})$. The rest follows as before. ∎

### v2

@@ 6. Free motivic sheaves
We already know that all truncations of $M\otimes\mathbb{Q}$ and $M\otimes\mathbb{Q}/\mathbb{Z}$ remain in $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$. To conclude, it suffices to control the transition map. Thus, take any $N,N^{\prime}\in\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$ that are concentrated in degree 0 and such that $N$ is rational while $N^{\prime}$ is torsion. We need to see that the image of any map $f:N\to N^{\prime}$ is in $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$. As this image is necessarily torsion, we have to see invariance under change of algebraically closed field. Thus let $C\subset C^{\prime}$ be an inclusion of non-discrete algebraically closed Banach fields over $X$. We want to see that the image of $N(C)\to N^{\prime}(C)$ maps isomorphically to the image of $N(C^{\prime})\to N^{\prime}(C^{\prime})$, noting that we know $N^{\prime}(C)=N^{\prime}(C^{\prime})$. Thus, the map is necessarily injective. But if a class $\alpha\in N^{\prime}(C)$ lies in the image of $N(C^{\prime})\to N^{\prime}(C^{\prime})=N^{\prime}(C)$, then by finitaryness there is some Banach-$C$-algebra $R$ of topologically finite type such that the image of $\alpha$ under $N^{\prime}(C)\to N^{\prime}(R)$ lies in the image of $N(R)\to N^{\prime}(R)$. But then there is a section $R\to C$, showing that $\alpha$ is already in the image of $N(C)$. ∎
@@ 7. The Cancellation Theorem
$\otimes\mathbb{Z}(1):\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)\to\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$
@@
We have to show that for all $M\in\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$, the map
$M\to\operatorname{\!{\mathscr{H}\!om}}(\mathbb{Z}(1),M\otimes\mathbb{Z}(1))$
is an isomorphism. The internal Hom here is taken in $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$, but it agrees with the internal Hom in the larger category $\mathcal{D}(X_{\mathrm{arc}},\mathbb{Z})$. Indeed, rewriting $\mathbb{Z}(1)$ in terms of $\mathbb{Z}_{\mathrm{mot}}[\mathbb{T}]/\mathbb{Z}$ for an annulus $\mathbb{T}$, and hence the internal Hom in terms of pushforward along $\mathbb{T}\to\ast$, this follows from Theorem 4.20. In particular, it now suffices to prove that for all strictly totally disconnected $X^{\prime}$ over $X$, the map
$M(X^{\prime})\to\mathrm{Hom}_{\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X^{\prime})}(\mathbb{Z}(1),M|_{X^{\prime}}\otimes\mathbb{Z}(1))$
is an isomorphism. Both sides here are naturally global sections of a sheaf on the profinite set $\mathcal{M}(X^{\prime})$, and as $\mathbb{Z}(1)$ is compact, we can even reduce to the case that $X^{\prime}$ (and then also $X$) is a geometric point.
Thus, we are reduced to the case of algebraically closed Banach fields $X=\mathcal{M}_{\mathrm{arc}}(C)$. The case of $C=\mathbb{C}$ is clear, so we can assume that $C$ is nonarchimedean. Also, the result is clear with torsion coefficients, so we will work with $\mathbb{Q}$-coefficients below.
@@
splitting the previous displayed map on $\pi_{i}$. Actually, we will rather have a map from the source to eventually defined sequences in the target, i.e. the colimit over all $N$ of $\mathrm{Map}(\mathbb{N}_{\geq N},-)$. Still, it will be a splitting in the sense that the composite map is the map taking an element to the constant sequence.
By Theorem 6.1, the left-hand side can be computed as the $i$-th homotopy group of the rationalization of the animated commutative monoid
@@
for some $k,\ell\geq 1$; and two such maps yield the same element if, possibly after refinement and addition of some section $\mathbb{T}\to\mathrm{Sym}^{k^{\prime}}(U\times\overline{\mathbb{G}}_{m})$, they admit a homotopy from $\mathbb{T}\times\mathrm{sd}^{\ell^{\prime}}\Delta^{i+1}_{C}\times\mathbb{B}$.
@@
This finishes the construction. It is easy to check that for all $N$ for which the map is defined, it is a splitting. Indeed, in that case $p_{1}=p_{2}$ and so $p_{1}^{N}=p_{2}$ cuts out the open balls around the N1N-1-th roots of unity (counted with multiplicity in case N1N-1 is divisible by the residue characteristic of $C$). The map to $\mathbb{N}[U]$ is constant, so $Y_{1}$ yields $N$ times the original section of $\mathbb{N}[U]$, while $Y_{2}$ yields it N1N-1 times; the difference is thus the original section. This already proves injectivity of the map
@@
We note that we are implicitly writing $\mathbb{Q}_{\mathrm{mot}}[U\times(\overline{\mathbb{G}}_{m},\ast)]$ as a retract of $\mathbb{Q}_{\mathrm{mot}}[U\times\overline{\mathbb{G}}_{m}]$, and we use that when we apply the above construction to maps to $\mathbb{Q}_{\mathrm{mot}}[U\times\ast]\subset\mathbb{Q}_{\mathrm{mot}}[U\times\overline{\mathbb{G}}_{m}]$, the construction yields zero. Indeed, in this case in the construction of $Y_{2}$ we have $p_{2}=1$ and hence we look at the locus $p_{1}^{N}=1$, which is precisely the locus of $N$-th roots of unity. Thus, $Y_{2}=Y_{1}$ in that case, and we get the zero map.
@@
${\operatorname{\!{\mathscr{H}\!om}}(\mathbb{Q}_{\mathrm{mot}}[(\mathbb{G}_{m},\ast)],\mathbb{Q}_{\mathrm{mot}}[U\times(\overline{\mathbb{G}}_{m},\ast)])\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{Q}_{\mathrm{mot}}[U]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{\!{\mathscr{H}\!om}}(\mathbb{Q}_{\mathrm{mot}}[(\mathbb{G}_{m},\ast)\times(\mathbb{G}_{m},\ast)],\mathbb{Q}_{\mathrm{mot}}[(\overline{\mathbb{G}}_{m},\ast)\times U\times(\overline{\mathbb{G}}_{m},\ast)])\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{\!{\mathscr{H}\!om}}(\mathbb{Q}_{\mathrm{mot}}[(\mathbb{G}_{m},\ast)],\mathbb{Q}_{\mathrm{mot}}[(\overline{\mathbb{G}}_{m},\ast)\times U])}$
@@
Of course, once this lemma was used to finish the proof of the cancellation theorem, we see that $\mathrm{End}(\mathbb{Z}(n))=\mathrm{End}(\mathbb{Z})$ which is concentrated in degree 0 and hence the triviality of the action of each $\sigma$ implies that the action of $\Sigma_{n}$ on $\mathbb{Z}(n)$ is homotopy-coherently trivial. For the moment, the weaker assertion is sufficient.
@@
There is an action of $\mathrm{GL}_{2}$ on $\mathbb{A}^{2}\setminus\{0\}$. We want to see that the antidiagonal matrix, switching the two coordinates, acts trivially on the reduced homology. But the subgroup generated by unipotent matrices acts trivially, by homotopy-invariance. This subgroup is $\mathrm{SL}_{2}$. Up the action of $\mathrm{SL}_{2}$, the antidiagonal matrix is equivalent to the diagonal matrix with entries (1,1)(-1,1). It remains to see that multiplication by 1-1 (or any element of $\mathbb{G}_{m}$) acts trivially on the reduced homology of $\mathbb{G}_{m}$, which is a direct verification using the identification with $\overline{\mathbb{G}}_{m}$. Indeed, the reduced homology of $\mathbb{G}_{m}$ is the kernel of $\mathbb{Z}_{\mathrm{mot}}[\mathbb{G}_{m}]\to\mathbb{Z}_{\mathrm{mot}}[\ast]$, and the isomorphism with $\overline{\mathbb{G}}_{m}$ is given by sending $\sum n_{i}[x_{i}]$ to $\prod\overline{x}_{i}^{n_{i}}$. The action of 1-1 sends $\sum n_{i}[x_{i}]$ to $\sum n_{i}[-x_{i}]$ which is mapped to $(-1)^{\sum n_{i}}\prod\overline{x}_{i}^{n_{i}}$. But $\sum n_{i}=0$ for any element in the kernel of $\mathbb{Z}_{\mathrm{mot}}[\mathbb{G}_{m}]\to\mathbb{Z}_{\mathrm{mot}}[\ast]$ and so $(-1)^{\sum n_{i}}=1$. ∎
@@ 8. $K$-theory
In the next corollary, we use that free motivic sheaves also exist for smooth Artin stacks, by left Kan extension.
@@
Of course, these results are also true with $\mathbb{Z}$-coefficients. We state it with $\mathbb{Q}$-coefficients here as our method of producing the splitting works only rationally.
@@
can be shown to be $\mathbb{Z}(n)[2n]$ as follows. Use the covering of $\mathbb{P}^{n}$ by $\mathbb{A}^{n}$ and $\mathbb{P}^{n}\setminus\{0\}$; the first piece is contractible, the second piece is an $\mathbb{A}^{1}$-fibration over $\mathbb{P}^{n-1}$ and hence cancels with $\mathbb{P}^{n-1}$, so the remainder is a shift of the reduced homology of the intersection $\mathbb{A}^{n}\setminus\{0\}$, which gives $\mathbb{Z}(n)[2n]$.) ∎
@@
We recall that $K_{\mathrm{cn}}(A)$ is (for us, by definition) the group completion of the $E_{\infty}$-monoid of finite projective $A$-modules. This yields an $E_{\infty}$-group, or equivalently a connective spectrum, cf. e.g. the appendix of [BS17], especially [BS17, Definition 12.13], for this perspective on $K$-theory.
We note that we are departing from our usual conventions, and consider a functor with values in spectra, not $\mathcal{D}(\mathbb{Z})$. As we will be working rationally momentarily, the difference will disappear. Note however that everything we did with $\mathcal{D}(\mathbb{Z})$-coefficients works also with $\mathrm{Sp}$-coefficients.
@@
Identifying connective spectra and $E_{\infty}$-groups in anima, $\mathbb{S}[\ast/\mathbb{G}_{m}]$ is the group completion of
@@
indeed, the displayed object is the free $E_{\infty}$-monoid on $\ast/\mathbb{G}_{m}$. Explicitly, this can be described as the $E_{\infty}$-monoid of finite projective modules $P$ equipped with lines $(L_{i})_{i}$, $L_{i}\subset P$, such that $\bigoplus_{i}L_{i}\to P$ is an isomorphism. We may then compute the rationalization of the group completion in terms of the Breen–Deligne resolution involving tensor products of terms
@@
embedding the normalizer of the diagonal torus (indeed, on the level of $E_{\infty}$-monoids, this takes $(P,(L_{i})_{i})$ to $P$), and it suffices to see that they induce isomorphisms
@@
We want to make $\overline{K}$ more explicit. For this, we first have to replace $K_{\mathrm{cn}}$ by a finitary approximation. We assume that our base ring $A$ is equipped with a unit $T\in A$ with $|T|<1$ and such that $T$ admits $n$-th roots for infinitely many $n$. In that case the non-unital ring $A_{<1}\subset A$ is Tor-unital: If $A_{<1}$ is an ideal in a unital ring $B$, then it is automatically the colimit of $B$ along multiplication by $T^{1/n_{i}-1/n_{i+1}}$ (where $n_{i}$ is a strictly increasing sequence of integers for which $T$ admits $n_{i}$-th roots) and hence flat and idempotent. This implies that the fiber of $K(B)\to K(B/A_{<1})$ is independent of $B$. Classically, this is due to Suslin [Sus95]. But even better, using Efimov’s $K$-theory of dualizable categories, we note that there is the Verdier quotient
$\mathcal{D}(B)\to\mathcal{D}(B/A_{<1})$
as $B/A_{<1}$ is an idempotent $B$-algebra. The kernel of this Verdier quotient is the dualizable $\infty$-category $\mathcal{D}(B^{a})$ of almost $B$-modules with respect to the ideal $A_{<1}$. As $K$-theory takes Verdier sequences to fibre sequences, we see that
$K(\mathcal{D}(B^{a}))=\mathrm{fib}(K(B)\to K(B/A_{<1})).$
But even $\mathcal{D}(B^{a})$ depends only on $A_{<1}$: Indeed, this category of almost modules is also the category of modules $M$ over the non-unital ring $A_{<1}$ such that the natural map
$M\otimes^{\mathbb{L}}_{A_{<1}}A_{<1}\to M$
is an isomorphism. Here, $M\otimes^{\mathbb{L}}_{A_{<1}}A_{<1}$ can be computed as the colimit of copies of $M$ with transition maps $T^{1/n_{i}-1/n_{i+1}}$ as above. We denote this $K$-theory by $K(A_{<1})$. Let $K_{\mathrm{cn}}(A_{<1})$ denote its connective cover. There is a natural map $K_{\mathrm{cn}}(A_{<1})\to K_{\mathrm{cn}}(A)$.
@@
The functor $A\mapsto\mathrm{cofib}(K_{\mathrm{cn}}(A_{<1})\to K_{\mathrm{cn}}(A))$, as well as its nonconnective version $\mathrm{cofib}(K(A_{<1})\to K(A))$, commutes with filtered colimits of Banach rings admitting a topologically nilpotent unit $T$ with roots as above. When evaluated on totally disconnected $A$, the connective and nonconnective versions agree.
@@
as $A=(\mathbb{Z}\oplus A_{<1})[T^{-1}]$. Passing to the cofiber of $K(A_{<1})$ the middle term becomes simply $K(\mathbb{Z})$ which is certainly invariant under changing $A$; and the left-most term is unchanged under replacing $A_{<1}$ by its $T$-adic completion, by the results of Thomason–Trobaugh [TT90, Proposition 3.19]. This gives the result for nonconnective $K$-theory. The case of connective $K$-theory follows.
@@
Fix some set of primes $S$. On strictly totally disconnected Banach rings $A$ in which $|p|\geq 1$ everywhere on $\mathcal{M}(A)$ for all $p\not\in S$, the functor $A\mapsto\mathrm{cofib}(K(A_{<1})\to K(A))[1/S]$ defines a finitary arc-sheaf of spectra that is ball-invariant, and hence agrees with $\overline{K}[1/S]$.
@@
After inverting the residue characteristic $p$ of $C$, this is a direct consequence of the identification with $\overline{K}(C)$. But after $p$-completion, $K(\mathcal{O}_{C})$ and $K(C)$ agree (cf. [HN20, Lemma 1.3.7]), so this cofiber agrees with $K(k)$ after $p$-completion, and this is just $\mathbb{Z}_{p}$ as $k$ is a perfect field (cf. [Kra80, Corollaire 5.5]). This has a unique $E_{\infty}$-$K(C)$-algebra structure. To glue the two pieces, we need to see that that the map
$\mathrm{cofib}(K(C_{<1})\to K(C))[1/p]\to\mathrm{cofib}(K(C_{<1})\to K(C))^{\wedge}_{p}[1/p]\cong\mathbb{Q}_{p}$
admits the structure of a map of $E_{\infty}$-$K(C)$-algebras. But the map is simply the projection $K(C)\to\mathbb{Z}$ given by the rank (which is trivial on $K(C_{<1})$), composed with inverting $p$ and embedding into $\mathbb{Q}_{p}$. ∎
@@
We can work with the fiber instead. This is the $K$-theory of the dualizable stable $\infty$-category $\mathcal{D}_{\mathrm{tor}}(A_{\leq 1})^{a}$ of torsion almost modules over $A_{\leq 1}$. Indeed, there are two Verdier sequences
@@
But $\mathcal{D}_{\mathrm{tor}}(A_{\leq 1})^{a}$ localizes on $\mathcal{M}(A)$ by the usual properties of the structure sheaf of perfectoid algebras, which in particular yields a sheaf of idempotent algebras with values in $\mathcal{D}_{\mathrm{tor}}(A_{\leq 1})^{a}$ (when interpreting torsion modules as complete modules). In other words, for any closed subset $Z\subset\mathcal{M}(A)$, the algebra $\mathcal{O}^{+}(Z)^{a}$ is an idempotent complete $A_{\leq 1}^{a}$-algebra, giving a Verdier sequence
$0\to\mathcal{D}_{\mathrm{tor}}(A_{\leq 1}^{a}\ \mathrm{on}\ U)\to\mathcal{D}_{\mathrm{tor}}(A_{\leq 1})^{a}\to\mathcal{D}_{\mathrm{tor}}(\mathcal{O}^{+}(Z)^{a})\to 0,$
where $U$ is the complement of $Z$ (and the term $\mathcal{D}(\ldots\ \mathrm{on}\ U)$ is defined as the fibre). If $Z$ and $Z^{\prime}$ cover $\mathcal{M}(A)$, then there are compatible Verdier sequences
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{D}_{\mathrm{tor}}(A_{\leq 1}^{a}\ \mathrm{on}\ U)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{D}_{\mathrm{tor}}(A_{\leq 1})^{a}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{D}_{\mathrm{tor}}(\mathcal{O}^{+}(Z)^{a})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{D}_{\mathrm{tor}}(\mathcal{O}^{+}(Z^{\prime})^{a}\ \mathrm{on}\ U)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{D}_{\mathrm{tor}}(\mathcal{O}^{+}(Z^{\prime})^{a})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{D}_{\mathrm{tor}}(\mathcal{O}^{+}(Z\cap Z^{\prime})^{a})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
where the left vertical functor is an equivalence. Applying $K$-theory then shows the desired sheaf property. ∎
@@
On perfectoid Banach rings $A$ (with $|p|_{A}<1$) such that all sheaves on the compact Hausdorff space $\mathcal{M}(A)$ are Postnikov complete, the functor $A\mapsto\mathrm{cofib}(K(A_{<1})\to K(A))_{\mathbb{Q}}$ is a finitary ball-invariant hypercomplete arc-sheaf. In particular, on such $A$, we have
@@
The condition that all sheaves are Postnikov complete is satisfied, for example, if $\mathcal{M}(A)$ has finite covering dimension [Lur09, Proposition 7.2.1.10].
@@
The commutation with filtered colimits comes from Proposition 8.8 and does not require any assumptions on $A$ and the $A_{i}$ with $A=\mathrm{colim}_{i}A_{i}$.
Rationally, finite étale descent holds because of trace maps; it even gives pro-finite étale descent. This combines with the analytic descent we already have to give a hypercomplete arc-sheaf. Indeed, from Theorem 4.1 we can produce a hypercomplete arc-sheaf preserving the values at strictly totally disconnected spaces. We need to see that this leaves the value at $A$ unchanged. But using pro-finite étale descent, we know that the value at totally disconnected $A$ stays unchanged, and then one uses descent along a profinite cover of $\mathcal{M}(A)$ (from analytic descent and Postnikov completeness) to see that the value at $A$ is also unchanged, under the assumption on $A$.
Finally, we check invariance under perfectoid balls (which is equivalent to usual ball-invariance). We know explicitly the value on $C\langle T^{1/p^{\infty}}\rangle_{1}$ (as $\mathcal{M}(C\langle T^{1/p^{\infty}}\rangle_{1})$ is a tree on which all sheaves are Postnikov complete): This is the cofiber of
@@
In particular, restricting to $A$ such that all sheaves on the compact Hausdorff space $\mathcal{M}(A)$ are Postnikov complete, the functor
@@
For the first statement, we can use the finitary nature of the functor to do a noetherian approximation to assume that $A_{\leq 1}$ is noetherian. The idea now is to use formal models and admissible blowups to relate analytic descent to (pro-)cdh descent results. We can either cite directly Cisinski’s cdh-descent for $KH$, [Cis13]. Alternatively, in order to give an argument independent of the existing motivic formalism, it follows from pro-cdh descent [Mor16, Theorem 3.7] (for the equal characteristic 0 case that we really need) and [KST18, Theorem A] (in the general case) upon making things ball-invariant. Namely, pro-cdh descent implies that the pro-spectrum $\{\mathrm{cofib}(K(A_{<r})\to K(A))\}_{r}$ satisfies descent, but after making it ball-invariant, all terms agree. Here $K(A_{<r})$ is notation for the fiber of $K(A_{\leq 1})\to K(A_{\leq 1}/A_{<r})$. The rest follows as before. ∎